At time t=0 a 2150 kg rocket in outer space fires an engine that exerts an increasing force on it in the +x−direction. This force obeys the equation Fx=At2, where t is time, and has a magnitude of 889.66 N when t=1.25 s. What impulse does the engine exert on the rocket during the 2.9 s interval starting 2.00 s after the engine is fired?
Answer #1

Soluna The expression for impulse with limit t=2.005 to t= (200 +299) :t= 24.99 re given by : .. - 4ogs 4.95 . Ygs. J= 1 Folt*5= 569.4 NV152 4999 - (2.0339] → J = 2081104 N.S